# Does Your Reasoning Model Implicitly Know When to Stop Thinking?

### v1

@@ 1 Introduction
Reinforcement learning from verifiable rewards (RLVR) algorithms, such as GRPO (Shao et al., 2024; Yang et al., 2026) and GSPO (Zheng et al., 2025a), have played a pivotal role in enabling test-time scaling. This capability allows large reasoning models (LRMs) like o3 (OpenAI, 2025a) and DeepSeek-R1 (Guo et al., 2025) to “think longer”. Longer CoTs enable LRMs to explore intermediate steps in greater depth and reduce abrupt logical leaps, thereby achieving unprecedented performance on challenging reasoning benchmarks such as AIME (Art of Problem Solving, 2024), OlympiadBench (Chaoqun et al., 2024) and IMO (Luong et al., 2025).
@@ A.1 Stimulating Reasoning Capabilities through Reinforcement Learning
The introduction of OpenAI o1 (OpenAI, 2025b) marks a major advance in reasoning performance and the beginning of the LRM era, inspiring efforts to replicate such strong reasoning abilities(Zou et al., 2025; Chen et al., 2025c; He et al., 2025). DeepSeek-R1, for example, achieves comparable results using a simple rule-based reward with the group relative policy optimization (GRPO)(Shao et al., 2024) algorithm, and its open-source release has established RLVR (DeepSeek-AI, 2025; Kimi Team, 2025a; Gao et al., 2024; Lambert et al., 2025; Zeng et al., 2025; Wen et al., 2025a; Song et al., 2025) as an effective paradigm for improving LLM reasoning, and Yang et al. (2026) provides a principled theoretical analysis of its advantage estimation. This paradigm simplifies reward design by employing binary 0/1 rewards determined through rule-based correctness evaluation, eliminating the need for separate reward models as required in original GRPO (Shao et al., 2024; DeepSeek-AI, 2025; Schulman et al., 2017; Huang et al., 2026) implementations, thereby substantially reducing memory and computational overhead during RL training.

### v2

@@ 1 Introduction
Reinforcement learning from verifiable rewards (RLVR) algorithms, such as GRPO (Shao et al., 2024; Yang et al., 2026) and GSPO (Zheng et al., 2025a; Lu et al., 2026), have played a pivotal role in enabling test-time scaling. This capability allows large reasoning models (LRMs) like o3 (OpenAI, 2025a) and DeepSeek-R1 (Guo et al., 2025) to “think longer”. Longer CoTs enable LRMs to explore intermediate steps in greater depth and reduce abrupt logical leaps, thereby achieving unprecedented performance on challenging reasoning benchmarks such as AIME (Art of Problem Solving, 2024), OlympiadBench (Chaoqun et al., 2024) and IMO (Luong et al., 2025).
@@ A.1 Stimulating Reasoning Capabilities through Reinforcement Learning
The introduction of OpenAI o1 (OpenAI, 2025b) marks a major advance in reasoning performance and the beginning of the LRM era, inspiring efforts to replicate such strong reasoning abilities(Chen et al., 2026; Zou et al., 2025; Chen et al., 2025c; He et al., 2025). DeepSeek-R1, for example, achieves comparable results using a simple rule-based reward with the group relative policy optimization (GRPO)(Shao et al., 2024) algorithm, and its open-source release has established RLVR (DeepSeek-AI, 2025; Kimi Team, 2025a; Gao et al., 2024; Lambert et al., 2025; Zeng et al., 2025; Wen et al., 2025a; Song et al., 2025) as an effective paradigm for improving LLM reasoning, and Yang et al. (2026) provides a principled theoretical analysis of its advantage estimation. This paradigm simplifies reward design by employing binary 0/1 rewards determined through rule-based correctness evaluation, eliminating the need for separate reward models as required in original GRPO (Shao et al., 2024; DeepSeek-AI, 2025; Schulman et al., 2017; Huang et al., 2026) implementations, thereby substantially reducing memory and computational overhead during RL training.
@@
In addition to the methods discussed above, a wide range of advanced techniques have been proposed in recent years to address various challenges in representation learning, model optimization, and generative modeling. These include progress in interpretable representation learning (Li et al., 2025a), prompt-based structural modeling (Li et al., 2025c), diffusion-driven restoration (Li et al., 2025b), efficient transformer architectures for visual modeling (Fu et al., 2022), prompt-guided sequence modeling (Cai et al., 2023, 2024), parameter-efficient tuning strategies (Cai et al., 2025a), as well as novel normalization mechanisms for improving model stability (Cai et al., 2025b). Although these works are designed for different task scenarios, they collectively enrich the toolkit of modern machine learning research and provide useful insights for understanding the generalization and optimization of neural models.
@@ D.4 Time Complexity Analysis
Appendix E Pseudo Code of SAGE
The example Python code of SAGE is illustrated in Figure 15, Figure 16 and Figure 17.
# Initialization and Parameter Processing
class SAGE:
[…]
self.llm = llm
self.tokenizer = tokenizer
self._init_special_tokens(eostep_token)
def forward(self, prompts, params):
# Process and validate input parameters
exploration_params = self._process_params(params)
# Initialize exploration instances for each prompt
instances = self._init_exploration_instances(prompts, exploration_params)
# Perform step-wise exploration reasoning
self._stepwise_exploration(instances, exploration_params)
# Generate final answers from completed reasoning
return self._generate_final_answers(instances, exploration_params)
def _init_special_tokens(self, eostep_token):
self.eot_token_id = self.tokenizer.convert_tokens_to_ids(’</think>’)
self.sot_token_id = self.tokenizer.convert_tokens_to_ids(’<think>’)
self.eostep_token_id = self.tokenizer.convert_tokens_to_ids(eostep_token)
self.eostep_token_list = [self.eostep_token_id, self.eot_token_id]
def _process_params(self, params):
processed = {}
processed[’exploration_width’] = params.get(’exploration_width’, 4)
processed[’max_ans_tokens’] = params.get(’max_ans_tokens’, 100)
processed[’max_step_num’] = params.get(’max_step_num’, 10)
processed[’per_step_token’] = params.get(’per_step_token’, 100)
processed[’enable_sort’] = params.get(’enable_sort’, True)
processed[’required_completions’] = params.get(’required_completions’, 1)
processed[’insert_sot’] = params.get(’insert_sot’, True)
processed[’top_p’] = params.get(’top_p’, 1.0)
processed[’temperature’] = params.get(’temperature’, 1.0)
return processed
def _init_exploration_instances(self, prompts, params):
instances = []
for prompt in prompts:
# Convert prompt to tokens if it’s a string
prompt_tokens = self.tokenizer.encode(prompt) if isinstance(prompt, str) else prompt
# Add start-of-thinking token if requested
if params[’insert_sot’]:
prompt_tokens.append(self.sot_token_id)
# Create exploration instance with empty candidates
instance = ExplorationInstance(
prompt_tokens,
logprobs=None,
required_completions=params[’required_completions’]
instances.append(instance)
return instances
Figure 15: Example Python code illustrating the SAGE Initialization and Parameter Processing.
# Step-wise Exploration
def _stepwise_exploration(self, instances, params):
for t in range(params[’max_step_num’]):
active_instances = [inst for inst in instances if not inst.is_finished]
if not active_instances:
break
for instance in active_instances:
# Initialize candidates if empty
if not instance.candidates:
self._init_instance_candidates(instance)
# Generate next reasoning steps
new_candidates = self._generate_next_steps(instance, params)
# Process and update candidates
self._update_instance_candidates(instance, new_candidates, params)
# Check if instance has enough completions
if len(instance.completed) >= instance.required_completions:
instance.is_finished = True
instance.candidates = []
def _init_instance_candidates(self, instance):
instance.candidates.append(ExplorationSequence(
tokens=instance.prompt_tokens,
logprobs=[],
cum_logprob=0.0))
def _generate_next_steps(self, instance, params):
# Create batch of prompts from current candidates
candidate_prompts = [cand.tokens for cand in instance.candidates]
# Configure step sampling parameters
step_params = SamplingParams(
n=params[’exploration_width’] * 2,
temperature=params[’temperature’],
top_p=params[’top_p’],
max_tokens=params[’per_step_token’],
stop_token_ids=self.eostep_token_list)
# Generate next steps using LLM
outputs = self.llm.generate(
prompt_token_ids=candidate_prompts,
sampling_params=step_params,
use_tqdm=False)
# Process outputs into new candidates
return self._process_step_outputs(instance.candidates, outputs)
def _process_step_outputs(self, current_candidates, outputs):
new_candidates = []
for i, (candidate, output) in enumerate(zip(current_candidates, outputs)):
if output and output.outputs:
for step_output in output.outputs:
new_candidate = ExplorationSequence(
tokens=candidate.tokens + step_output.token_ids,
logprobs=candidate.logprobs + step_output.logprobs,
cum_logprob=candidate.cum_logprob + step_output.cumulative_logprob)
new_candidates.append(new_candidate)
return new_candidates
Figure 16: Example Python code illustrating the Step-wise Exploration of SAGE.
# Candidate Management and Scoring
def _update_instance_candidates(self, instance, new_candidates, params):
completed, ongoing = self._separate_completed_candidates(new_candidates)
instance.completed.extend(completed)
if params[’enable_sort’]:
ongoing.sort(key=self._score_candidate, reverse=True)
instance.candidates = ongoing[:params[’exploration_width’]]
def _separate_completed_candidates(self, candidates):
completed = []
ongoing = []
for candidate in candidates:
if candidate.tokens[-1] == self.eot_token_id:
completed.append(candidate)
else:
ongoing.append(candidate)
return completed, ongoing
def _score_candidate(self, candidate):
return candidate.cum_logprob / len(candidate.tokens)
# Final Answer Generation
def _generate_final_answers(self, instances, params):
outputs = []
for instance in instances:
# Combine and sort all candidates
all_candidates = instance.completed + instance.candidates
all_candidates.sort(key=self._score_candidate, reverse=True)
# Get best candidates and ensure proper formatting
best_candidates = all_candidates[:instance.required_completions]
self._finalize_candidates(best_candidates)
# Generate answers based on completed reasoning
answer_candidates = self._generate_answers_from_candidates(best_candidates, params)
# Create output structure
outputs.append(ExplorationOutput(sequences=answer_candidates))
return outputs
def _finalize_candidates(self, candidates):
for candidate in candidates:
if self.eot_token_id not in candidate.tokens:
candidate.tokens.append(self.eot_token_id)
def _generate_answers_from_candidates(self, candidates, params):
candidate_prompts = [cand.tokens for cand in candidates]
answer_params = SamplingParams(
n=1, temperature=0.0,
max_tokens=params[’max_ans_tokens’])
answer_outputs = self.llm.generate(
prompt_token_ids=candidate_prompts,
sampling_params=answer_params,use_tqdm=False)
for candidate, output in zip(candidates, answer_outputs):
if output and output.outputs:
candidate.tokens += output.outputs[0].token_ids
candidate.text = self.tokenizer.decode(candidate.tokens)
return candidates
Figure 17: Example Python code illustrating the Candidate Management, Scoring and Final Answer Generation of SAGE.